OpenStudy (anonymous):

1. (a) Calculate the formula mass of Fe2(SO4)3. (b) What is the molar mass of Fe2(SO4)3? (c) What is the mass of 0.2000 mole of Fe2(SO4)3? 2. Calculate the number of atoms in 27.0 g of pure silver.

OpenStudy (anonymous):

Iron: 55.85g /mole Sulfur: 32.06g/mole Oxygen: 16.00g/mole I can't recall but I think a and b are the same? So 55.85x2=111.7 g /mol 32.06x3=96.18 g /mol 16.00x12=192.0 g / mol 399.9 g /mol c) 0.2000 mol x 399.9 g /mol = 5.001x10^-4

OpenStudy (anonymous):

I also think that molar mass and formula mass are the same.... I am confused in part c. Don't need to have 27.0 in the equation?

OpenStudy (anonymous):

isn't 0.2000 mol x 399.9 g /mol = 79.98 ??

OpenStudy (anonymous):

No, I think those are separate questions. Number 2: 27.0g Ag x 1 mol / 107.87 g Ag x 6.022x10^23 / 1 mol = 1.51 x 10^23 atoms Ag

OpenStudy (anonymous):

And yes, I think I must be very tired, lol.. Sorry!

OpenStudy (anonymous):

never mind. thank you

OpenStudy (anonymous):

would you kindly help me with another question?

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

Calculate the mass of (a) 0.200 mol of chlorine atoms. (b) 0.200 mol of chlorine molecules. (c) chlorine which contains the same number of molecules as there are in 1.20 mol of water. I am confused in masses and mole, what is the formulas about masses and mole?

OpenStudy (anonymous):

Ok, if you are given moles and it asks for atoms, then you use Avogadro's number. If you are given moles and it asks for molecules, you still use Avogadro's number. So a and b would be the same for this one.

OpenStudy (anonymous):

Do you know my question? Biochem? lol

OpenStudy (anonymous):

I am not studying Bio but Chem, sorry that I cannot understand the question about bio...

OpenStudy (anonymous):

lol me either...

OpenStudy (anonymous):

atoms and molecules are not the same right?

OpenStudy (anonymous):

well no, but Avogadro's number is used to calculate both. http://science.howstuffworks.com/avogadros-number.htm

OpenStudy (anonymous):

I mean the definition of atoms and molecules.

OpenStudy (anonymous):

Well a molecule is made up of multiple atoms

OpenStudy (anonymous):

oh I almost forgot it .......

OpenStudy (anonymous):

the mass of 0.200 mol of chlorine atoms the moles is 0.200 right?

OpenStudy (anonymous):

the number of mole is 0.200?

OpenStudy (anonymous):

the question above was the number of molecules/atoms in 0.200 moles of chlorine

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

there are 0.200 mol of chlorine atoms and I need to find out the mass right?

OpenStudy (anonymous):

simply convert from moles to atoms... .200 mol x 6.022x10^23/mol = 1.20 atoms/molecules of chlorine...You do NOT need to find the mass for this...

OpenStudy (anonymous):

6.02*10^(23)*0.200=1.204^(23) right

OpenStudy (anonymous):

ummm...

OpenStudy (anonymous):

yep that works too

OpenStudy (anonymous):

Remember sic figs

OpenStudy (anonymous):

but the question is asking for the mass....isn't it?

OpenStudy (anonymous):

I don't really know how to say this so you'll get it lol...No.

OpenStudy (anonymous):

...

OpenStudy (anonymous):

part b is the same as part a?